Given f(x)=x+1 and h(x)= 2x-3, find (f-h)(x).
Found 2 solutions by addingup, MathTherapy:
Answer by addingup(3677)

If:
f(x) = x + 1
and
h(x) = 2x - 3
then
f - h = x + 1 - 2x - 3
group like terms:
f - h = x -2x + 1 -3
simplify by subtraction to get your answer

Answer by MathTherapy(10552)

Given f(x)=x+1 and h(x)= 2x-3, find (f-h)(x).
f(x) = x + 1          h(x) = 2x - 3
Correct answer:
IGNORE anyone who says otherwise, as he/she is leading you astray!